OpenStudy (anonymous):

Shalika bought a purse for $120. The tax rate is 9%. Which proportion can be used to find the amount of tax Shalika will pay? A. 9/120=x/100 B. 120/x=100/9 C. x/120=100/9 D. 120/9=100/x

OpenStudy (paki):

have you any idea here...?

OpenStudy (anonymous):

nope

OpenStudy (karatechopper):

ok, so 9%, how can you write that in decimal form?

OpenStudy (karatechopper):

or actually, in fraction form.

OpenStudy (anonymous):

0.9

OpenStudy (anonymous):

?

OpenStudy (skullpatrol):

B. 120/x=100/9

OpenStudy (skullpatrol):

Any questions @Roc07or ?

OpenStudy (skullpatrol):

@Roc07or are you still there?

OpenStudy (skullpatrol):

Try to think about the defining relationship for percents: $$\Huge \frac{x}{100}=\frac{a}{b}$$

OpenStudy (skullpatrol):